# Multi-Stage Active Sequential Hypothesis Testing with Clustered Hypotheses

### v1

@@ III-A Preprocessing Step - Per-Action Hypothesis Clustering
In this subsection, we explain the hypothesis clustering concept in depth. The main purpose of the hypothesis clustering is to allow the DM to refine its “bad actions,” characterized by small KLD between some of their output distributions (e.g., either zero or very close to $\alpha$ from assumption (A1)). Namely, to tell apart between two or more very close hypotheses, the DM would require, on average, a large number of observations. Therefore, since these hypotheses are very similar, the DM can significantly reduce the number of the collected observations by simply identifying the set of similar hypotheses (e.g., by identifying a single member of these hypotheses as we do in this work) and using a different action with better separation on the found set in the next stage. In other words, the DM can improve its performance in terms of mean detection delay (and, consequently, obtain lower ABR) by clustering hypotheses by their informational proximity defined by a proximity parameter $\varepsilon$ set by the user.
Hypothesis clustering by their proximity (i.e., selecting $\varepsilon$) and selecting a representative from each cluster are not trivial. Thus, we first discuss a key property of a good hypothesis clustering mechanism. Assume action $a$ is repeatedly used and that $H_{i}$ is the underlying hypothesis. Now, we observe the normalized accumulated LLR between $H_{k}$ in $H_{i}$’s cluster against $H_{j}$ that is not in $H_{i}$’s cluster converges to
$\frac{1}{n}L_{kj}$ $\to\mathbb{E}\left[\log\frac{f_{a}(X;{\underline{\theta}}_{k}(a))}{f_{a}(X;{\underline{\theta}}_{j}(a))}\middle|H_{i}\right]$
$=\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{i}(a))\|f_{a}(\cdot;{\underline{\theta}}_{j}(a)))$
$\quad\quad-\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{i}(a))\|f_{a}(\cdot;{\underline{\theta}}_{k}(a)))\triangleq\Delta\mathcal{D}_{ijk}$ (2)
in probability from the Asymptotic Equipartition Property (AEP) property [21, Theorem 11.8.1]. Since $\frac{1}{n}\gamma\to 0$ for any fixed $\gamma$, we observe that $H_{i}$ can be declared as true if and only if the difference in (2) is positive. Therefore, any $H_{k}$ clustered with $H_{i}$ should have its KLD (from $H_{i}$ to $H_{k}$) strictly lower than $H_{i}$’s KLD from any $H_{j}$ not clustered with $H_{i}$.
@@
Now, we can formalize the desired clustering on the hypotheses using their parameter space. Let $\operatorname{param}\left(a\right)\triangleq\{{\underline{\theta}}_{i}(a):i\in\mathcal{H}\}$ be the set of distribution parameters under action $a$. Let $\mathcal{B}_{\varepsilon}({\underline{\nu}})\triangleq\{{\underline{\theta}}\in\operatorname{param}\left(a\right):\|{\underline{\theta}}-{\underline{\nu}}\|_{2}^{2}\leq\varepsilon\}$ be the $\varepsilon$-neighborhood of ${\underline{\nu}}$ composed of points from $\operatorname{param}\left(a\right)$. We say that ${\underline{y}}$ is reachable from ${\underline{x}}$ if there are ${\underline{x}}_{1},{\underline{x}}_{2},\dots,{\underline{x}}_{r-1}$ such that ${\underline{x}}={\underline{x}}_{1}$, ${\underline{x}}_{j}\in\mathcal{B}_{\varepsilon}({\underline{x}}_{j+1})$ and, additionally we have ${\underline{x}}_{r-1}\in\mathcal{B}_{\varepsilon}({\underline{y}})$. Similarly, ${\underline{x}}$ and ${\underline{y}}$ are connected if ${\underline{x}}$ is reachable from ${\underline{y}}$ and vice-versa.
Thus, we define the cluster of $H_{i}$ as $\mathcal{C}_{i}(a,\varepsilon)\subseteq\operatorname{param}\left(a\right)$ to be a non-empty subset such that (i) for any ${\underline{\nu}}$, if ${\underline{\nu}}$ is reachable from ${\underline{\theta}}_{i}$ then ${\underline{\nu}}\in\mathcal{C}_{i}(a,\varepsilon)$, and (ii) any ${\underline{\nu}},{\underline{\theta}}\in\mathcal{C}_{i}(a,\varepsilon)$ are connected. Now, we define
$\operatorname{equiv}\left(i,a\right)$$\triangleq\left\{k\in\mathcal{H}\ :\ {\underline{\theta}}_{k}\in\mathcal{C}_{i}(a,\varepsilon)\right\}$
@@
to be the set of equivalent hypotheses to $H_{i}$ and the set of representatives taken from $\mathcal{U}\subseteq\mathcal{H}$, respectively. Note that the choice of the smallest indices is arbitrary. If no clustering is used, we define $\operatorname{equiv}\left(i,a\right)\triangleq\{j\in\mathcal{H}:{\underline{\theta}}_{j}(a)={\underline{\theta}}_{i}(a)\}$ instead. Remark: We assume that even after the clustering, assumption (A3) holds.
The use of the $\ell_{2}$-norm allows us to use rich density-based clustering over $\mathbb{R}^{M_{a}}$ (e.g., [22, Chapter 5]) to cluster hypotheses by proximity. Notably, the famous Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [23] is tailored to our needs and is sufficient for clustering. DBSCAN groups together points with many shared neighbors (specified by the MinPts parameter) within their $\varepsilon$-neighborhood and labels all remaining points as noise or outliers. Furthermore, DBSCAN guarantees that clusters are $\varepsilon$ apart from each other.
@@ III-B Multi-Staged Algorithm
In this subsection, we present and describe our multi-stage algorithm. The procedure is given in Algorithm 1. In Line 1, we initialize the “alive” hypotheses, $H_{\mathrm{alive}}$, to be the set of all hypotheses. The algorithm thins this set in each stage until one or no hypothesis remains. If a single hypothesis remains at the end of the procedure, it is declared as true, and an error is raised otherwise.
We term Lines 3-11 as stages. In each stage, an action $a$ is computed (Line 3). Specifically, each action is associated with a non-zero minimal distance between each two alive hypotheses, and the action with the maximal minimum distance is selected for the current round. Line 4 computes the representatives of each cluster out of the alive hypotheses. Now, we apply action $a$ repeatedly until a winner against all hypotheses, $i^{*}$, is found (Lines 6-9). Then, we update $H_{\mathrm{alive}}$ to be the alive hypotheses in $i^{*}$’s cluster (Lines 10-11).
@@ IV Analysis
In this section, we analyze Algorithm 1 in terms of error probability, mean detection delay, and the scaling of the ABR as functions of $\delta$ and $H$. Notably, all previous results regarding the Armitage Test, e.g., vanishing error probability or bounded mean detection delay, do not trivially follow from using the Armitage Test at each stage due to the clustering, and accordingly, all computations must be modified. Due to brevity, the proofs are attached in Appendix A.
We start with the error probability. To bound the error probability of the algorithm, we first bound the per-stage error probability. Each stage of the algorithm is terminated when some hypothesis has all of its LLRs exceed $\log\frac{H}{\delta}$. At first glance, there are two possible error events; the first is that the stage winner (at some stage) is not the representative of the real hypothesis, and the second is the event when two different hypotheses win in the current round. We argue that the latter case never occurs in the following lemma:
@@
See Appendix A-A. ∎
Now that we have established that there are no ties, we bound the former error probability with the following lemma:
Lemma 2.
For any action $a$, there exists some $\eta_{a}\in[0,1)$ such that $\Delta\mathcal{D}_{ijk}>\eta_{a}$, the current action still can distinguish between at least two hypotheses, and for any $\varepsilon\leq\eta_{a}$, the per-stage error probability cannot exceed $\frac{\delta}{H}$. Particularly, $p_{e}\leq\delta$.
Proof:
See Appendix A-B. ∎
Now, we focus on the scaling laws of the mean detection delay, $\mathbb{E}\left[N\right]$, given in the following lemma:
Lemma 3.
Let $\tau_{r}$ be the number of observations used in stage $r$ when using action $a_{r}$, $H_{\mathrm{alive}}^{(r)}$ be the alive hypotheses in stage $r$, and $k$ be the representative from $H_{i}$’s cluster. Then $\mathbb{E}\left[\tau_{r}|H_{i}\right]\approx\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ for any $i$. Furthermore, $\mathbb{E}\left[\tau_{r}|H_{i}\right]$ is bounded, i.e., $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[\tau_{r}|H_{i}\right]\leq\frac{1}{\varepsilon}\log\frac{H}{\delta}$ and particularly, $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\frac{H}{\varepsilon}\log\frac{H}{\delta}$.
Proof:
See Appendix A-C. ∎
Remark: Although $\mathbb{E}\left[N\right]$ seems identical to non-clustering cases, it is, in fact, smaller as the expectation not only contains fewer positive addends, but each addend is also smaller.
Corollary 1 (Asymptotic Optimality).
For a suitable choice of $\varepsilon$, Algorithm 1 has vanishing ABR when $\delta\to 0$. Furthermore, the ABR does not exceed $\delta$ when $H\to\infty$.
Proof:
See Appendix A-D. ∎
V Numerical Results
In this section, we present numerical results illustrating the performance of our algorithm compared to GJL. To this end, we have conducted simulations. Namely, we compare GJL to our algorithm in two scenarios: (i) when no clustering is used. (ii) clustering is used with a suitable proximity parameter.
In the simulations, we have used uniform priors, i.e., $\pi_{i}=\frac{1}{H}$. For a fair comparison, we picked distributions whose mean can distinguish between different hypotheses, namely, the exponential and unit-variance normal distributions. The number of hypotheses was set to $H=16$, and the DM had $|\mathcal{A}|=17$ actions. The first $H$ actions had a mean of 3 if $H_{a-1}$ is true for $1\leq a\leq H$, and a random mean drawn uniformly from $[0,1]$ otherwise. The last action had the means follow the equation $0.5+0.01\times i$ for $0\leq i\leq H$. All random means were drawn only once. For the clustering algorithm, DBSCAN, we use $\text{minPts}=1$ and proximity parameters $\varepsilon\in\{0,0.1\}$.
In Figure 2, we present the results for the two cases. The dashed lines present the ABR for the case where all actions produce observations following the normal distribution, whereas non-dashed lines present the ABR for the case where all actions produce exponentially distributed observations. The orange curves correspond to GJL, and the blue and green curves correspond to our algorithm without and with clustering, respectively. Each algorithm attains vanishing ABR as $\delta\to 0$, but the non-greedy selections made by each instance of our algorithm enjoys significantly lower ABR, which is attributed to lower mean detection delay.
[…]
Figure 2: Comparison of the ABR (equation (1)) for $H=16$ using our algorithm (with proximity parameter $\varepsilon\in\{0,0.1\}$) against GJL. The observations obtained are either normally distributed (dashed) or exponentially distributed (non-dashed). The ABR vanishes as $\delta\to 0$ for all instances, but GJL’s greedy selection has a slower decay.
Appendix A Miscellaneous Proofs
A-A Proof of Lemma 1
Assume to the contradiction that there are two winners, $i_{r}^{*}$ and $j_{r}^{*}$. Since $i_{r}^{*}$ is a winner, then $L_{i_{r}^{*}j_{r}^{*}}\geq\log\frac{H}{\delta}$. Likewise, $j_{r}^{*}$ is a winner so $L_{j_{r}^{*}i_{r}^{*}}\geq\log\frac{H}{\delta}$. However, due to symmetry, we have $\log\frac{H}{\delta}\leq L_{i_{r}^{*}j_{r}^{*}}=-L_{j_{r}^{*}i_{r}^{*}}\leq-\log\frac{H}{\delta}$, contradiction.
A-B Proof of Lemma 2
Since the number of alive hypotheses can only decrease in each stage, there can be no more than $H$ stages. Consequently, we can compute the error probability of a single stage, and then, by the union bound, the error probability of the algorithm, $p_{e}$, is at most $H$ times the error probability of a single stage. Hence, we will focus on computing the first stage’s error probability for simplicity.
@@ A-B Proof of Lemma 2
A-C Proof of Lemma 3
Much like in the proof of the previous lemma, we notice that the Armitage Test essentially conducts $\operatorname{\mathcal{O}}\left(H^{2}\right)$ Wald Tests in parallel using the same observation vector. Due to the algorithm’s stage termination condition when clustering, we are interested in $\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]$ and the number of observations required for it to exceed $\log\frac{H}{\delta}$, where $k$ is the index of the hypothesis representing $H_{i}$ like before. Using Wald’s Identity, we have
$\mathbb{E}\left[L_{kj}|H_{i}\right]$ $=\mathbb{E}\left[\sum_{t=1}^{\tau_{r}}\log\frac{f_{a_{r}}(x_{t};{\underline{\theta}}_{k}(a_{r}))}{f_{a_{r}}(x_{t};{\underline{\theta}}_{j}(a_{r}))}\middle|H_{i}\right]$
$=\mathbb{E}\left[\tau_{r}|H_{i}\right]\times\mathbb{E}\left[\log\frac{f_{a_{r}}(X;{\underline{\theta}}_{k}(a_{r}))}{f_{a_{r}}(X;{\underline{\theta}}_{j}(a_{r}))}\middle|H_{i}\right]$
$=\mathbb{E}\left[\tau_{r}|H_{i}\right]\times\Delta\mathcal{D}_{ijk}$ (5)
for any relevant $i$, $k$, and $j$ (and particularly, for the LLR minimizer). Denote $\hat{p}_{e}$ as the per-stage error probability. From the stopping rule, we have
$\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]=\hat{p}_{e}\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i},\exists j\neq i\text{ wins}\right]$
$+(1-\hat{p}_{e})\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i},\min_{l}L_{kl}\geq\log\frac{H}{\delta}\right]$
using Bayes. Now, in similar arguments made by Wald in [1, Chapter 4], we approximate $\hat{p}_{e}\approx\frac{\delta}{H}$ for small enough $\delta$. Furthermore, conditioned on success, we approximate $\min_{l}L_{kl}\approx\log\frac{H}{\delta}$ for small enough $\delta$. On the other hand, conditioned on error, we observe that $\min_{l}L_{kl}<\log\frac{H}{\delta}$ since otherwise there would be no error as the winner is unique per Lemma 1. This inequality can be weakened, and we obtain that
$\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]$ $\lesssim\left(1-\frac{\delta}{H}\right)\log\frac{H}{\delta}+\frac{\delta}{H}\log\frac{H}{\delta}$
$=\log\frac{H}{\delta}.$ (6)
Specifically, combining (5) and (6) yields
$\mathbb{E}\left[\tau_{r}|H_{i}\right]$ $\lesssim\frac{\log\frac{H}{\delta}}{\min_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\Delta\mathcal{D}_{ijk}}$
$=\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ (7)
and the upper bound follows. Finally, since $i$ and $j$ are not clustered together, $\Delta\mathcal{D}_{ijk}\geq\varepsilon$ and we bound (7) by $\frac{1}{\varepsilon}\log\frac{H}{\delta}$. The lower bound
$\mathbb{E}\left[\tau_{r}|H_{i}\right]$ $\geq\frac{\log\frac{H}{\delta}}{\min_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\Delta\mathcal{D}_{ijk}}$
$=\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$
$\geq\frac{\log\frac{H}{\delta}}{\beta}$
follows from applying [24, Theorem 3.1] (using the law of large numbers as in (2)) as long as $\Delta\mathcal{D}_{ijk}>0$ for any relevant $i$, $k$ and $j$. Notably, the set $j$ is taken from, $\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}$, is the set of possible hypotheses to participate in the Armitage Test against $H_{i}$’s representative. Finally, by averaging over the underlying hypothesis using the priors $\{\pi_{i}\}_{i=0}^{H-1}$, we obtain that
$\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[\tau_{r}\right]\leq\frac{1}{\varepsilon}\log\frac{H}{\delta}$
as desired. All that remains is bounding the number of stages; There is at least one stage and at most $H-1$ stages, so
$\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\frac{H}{\varepsilon}\log\frac{H}{\delta}.$
A-D Proof of Corollary 1
Combining Lemma 2 and Lemma 3 with (1) yields
$\frac{\delta}{H^{2}}\mathbb{E}\left[N\right]+p_{e}\leq\frac{\delta}{H^{2}}\times\frac{H}{\varepsilon}\log\frac{H}{\delta}+\delta=\frac{1}{\varepsilon}\times\frac{1}{\frac{H}{\delta}}\log\frac{H}{\delta}+\delta$
, which vanishes when $\delta\to 0$. When $H\to\infty$, only the first addend vanishes, and we are left with $\delta$.

### v4

@@ III-A Preprocessing Step - Per-Action Hypothesis Clustering
In this subsection, we explain the hypothesis clustering concept in depth. The main purpose of the hypothesis clustering is to allow the DM to refine its “bad actions,” characterized by small KLD between some of their output distributions (e.g., either zero or very close to $\alpha$ from Assumption (A1)). Namely, to distinguish between two or more very close hypotheses, the DM would require, on average, a large number of samples. Therefore, since these hypotheses are very similar, the DM can significantly reduce the number of the collected samples by simply identifying the set of similar hypotheses (e.g., by identifying a single member of these hypotheses as we do in this work) and using a different action with better separation on the found set in the next stage. In other words, the DM can improve its performance in terms of mean detection delay (and, consequently, obtain lower ABR) by clustering hypotheses by their informational proximity defined by a proximity parameter $\varepsilon_{a}$ set by the user.
Clustering hypotheses by their proximity (i.e., selecting $\varepsilon_{a}$) and selecting a representative from each cluster is not trivial. Thus, we first discuss a key property of a good hypothesis clustering mechanism. Assume action $a$ is repeatedly used and that $H_{i}$ is the underlying hypothesis. Now, we observe the normalized accumulated LLR between $H_{k}$ in $H_{i}$’s cluster against $H_{j}$ that is not in $H_{i}$’s cluster satisfies
$\frac{1}{n}L_{kj}$ $\xrightarrow{n\to\infty}\mathbb{E}\left[\log\frac{f_{a}(X;{\underline{\theta}}_{k}(a))}{f_{a}(X;{\underline{\theta}}_{j}(a))}\middle|H_{i}\right]\triangleq\Delta\mathcal{D}_{ijk}(a)$
$=\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))-\mathcal{D}_{KL}(H_{i}(a)\|H_{k}(a))$ (2)
by the Asymptotic Equipartition Property (AEP) [22, Theorem 16.8.1], when the convergence is almost sure convergence. Like before, we will drop $a$ when it is clear from context. If the difference in (2) is positive, then $H_{k}$ would win against $H_{j}$ when a sufficiently large number of samples is collected when $H_{i}$ is true. Therefore, any $H_{k}$ clustered with $H_{i}$ should have $\mathcal{D}_{KL}(H_{i}(a)\|H_{k}(a))<\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))$ for any $H_{j}$ not clustered with $H_{i}$.
@@
Now, we can formalize the desired clustering on the hypotheses using their parameter space. Let $\operatorname{param}\left(a\right)\triangleq\{{\underline{\theta}}_{i}(a):i\in\mathcal{H}\}$ be the set of distribution parameters under action $a$. Let $\mathcal{B}_{\varepsilon_{a}}({\underline{\nu}})\triangleq\{{\underline{\theta}}\in\operatorname{param}\left(a\right):\|{\underline{\theta}}-{\underline{\nu}}\|_{2}^{2}\leq\varepsilon_{a}\}$ be the $\varepsilon_{a}$-neighborhood of ${\underline{\nu}}$ composed of points from $\operatorname{param}\left(a\right)$. We say that ${\underline{y}}$ is reachable from ${\underline{x}}$ if there are ${\underline{x}}_{1},{\underline{x}}_{2},\dots,{\underline{x}}_{r-1}$ such that ${\underline{x}}={\underline{x}}_{1}$, ${\underline{x}}_{j}\in\mathcal{B}_{\varepsilon_{a}}({\underline{x}}_{j+1})$ (for $21\leq j\leq r-$ 2) and, additionally we have ${\underline{x}}_{r-1}\in\mathcal{B}_{\varepsilon_{a}}({\underline{y}})$.
Thus, $H_{i}$’s cluster, $\mathcal{C}_{i}(a,\varepsilon_{a})\subseteq\operatorname{param}\left(a\right)$, is a non-empty subset such that for any ${\underline{\nu}}$, if ${\underline{\nu}}$ is reachable from ${\underline{\theta}}_{i}$ then ${\underline{\nu}}\in\mathcal{C}_{i}(a,\varepsilon_{a})$. Now, we define
$\operatorname{equiv}\left(i,a\right)$$\triangleq\left\{k\in\mathcal{H}\ :\ {\underline{\theta}}_{k}\in\mathcal{C}_{i}(a,\varepsilon_{a})\right\}$
@@
to be the set of equivalent hypotheses to $H_{i}$ and the set of representatives taken from $\mathcal{U}\subseteq\mathcal{H}$, respectively. $\operatorname{repr}\left(\mathcal{U},a\right)$ will contain the set of contestants (out of $\mathcal{U}$) in each stage of our strategy. Note that the choice of the smallest indices is arbitrary, and that Assumption (A3) still holds (for $\mathcal{U}$) if $|\operatorname{repr}\left(\mathcal{U},a\right)|>1$. If no clustering is used, we define $\operatorname{equiv}\left(i,a\right)\triangleq\{j\in\mathcal{H}:{\underline{\theta}}_{j}(a)={\underline{\theta}}_{i}(a)\}$ instead.
The use of the $\ell_{2}$-norm allows us to use rich density-based clustering over $\mathbb{R}^{M_{a}}$ (e.g., [23, Chapter 5]) to cluster hypotheses by proximity. Notably, the famous Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [24] is tailored to our needs and is sufficient for clustering. DBSCAN groups together points with many shared neighbors (specified by the MinPts parameter) within their $\varepsilon_{a}$-neighborhood and labels all remaining points as noise or outliers. Furthermore, DBSCAN guarantees that clusters are $\varepsilon_{a}$ apart from each other.
@@ III-B Multi-Staged Algorithm
In this subsection, we present and describe our multi-stage algorithm. The procedure is given in Algorithm 1. In Line 1, we initialize the “alive” hypotheses, $H_{\mathrm{alive}}$, to be the set of all hypotheses. The algorithm thins this set in each stage until one or no hypotheses remain. If a single hypothesis remains at the end of the procedure, it is declared true.
We term Lines 3-11 as stages. At the beginning of each stage, the action $a$ with the largest minimum distance between the alive hypotheses (i.e., best separation) is selected (Line 3). Line 4 computes the cluster representatives out of the alive hypotheses. This set contains the contestants for the current stage. Action $a$ is repeatedly applied until a winner against all hypotheses, $i^{*}$, is found (Lines 6-9). Then, $H_{\mathrm{alive}}$ is updated to be the alive hypotheses in $i^{*}$’s cluster (Lines 10-11).
@@ IV Analysis
In this section, we analyze Algorithm 1 in terms of error probability, mean detection delay, and the scaling of the ABR as functions of $\delta$ and $H$. Notably, previous results regarding the Armitage Test, e.g., vanishing error probability, do not trivially follow when clustering hypotheses, and accordingly, some computations must be modified. When no clustering is used, however, the previous results naturally follow. We start with stating the main result:
Theorem 1.
Let $\tau_{r}$ be the number of samples used in stage $r$ when using action $a_{r}$, $H_{\mathrm{alive}}^{(r)}$ be the alive hypotheses in stage $r$, and $k$ be the representative from $H_{i}$’s cluster. Then $\mathbb{E}\left[\tau_{r}|H_{i}\right]=(1+o(1))\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ for any $i$ where the little-O is with respect to $\delta\to 0$. Furthermore, $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\max_{a\in\mathcal{A}}\left\{\frac{H}{\varepsilon_{a}}\log\frac{H}{\delta}\right\}$.
Proof:
The result on $\mathbb{E}\left[\tau_{r}|H_{i}\right]$ is a direct result of applying [4, Theorem 4.1] over the set of hypotheses participating in stage $r$’s Armitage Test. Notably, the theorem is applicable only if $\Delta\mathcal{D}_{ijk}>0$, i.e., for an appropriate choice of $\varepsilon_{a}$. The bounds on $\mathbb{E}\left[N\right]$ follow since the number of stages is at least one and at most $H-1$. ∎
We argue that the non-hypothesis clustering flavor of our algorithm improves over GJL if both use the same action sequence. Let $\mathcal{GJL}(a,\mathcal{U})=\{(l,m)\in\mathcal{U}^{2}:\mathcal{D}_{KL}(H_{l}(a)\|H_{m}(a))>0\}$ be the set of pairs from a hypotheses subset $\mathcal{U}\subseteq\mathcal{H}$ whose KLD under action $a$ is not zero. Let $\mathcal{D}_{\min}(a,\mathcal{U})=\mathop{\operatorname{argmin}}_{(l,m)\in\mathcal{GJL}(a,\mathcal{U})}\left\{\mathcal{D}_{KL}(H_{l}(a)\|H_{m}(a))\right\}$ be the smallest non-zero KLD of hypotheses $\mathcal{U}$ under action $a$. Regardless of which hypothesis is true, if GJL applies action $a$, it uses $\tau_{GJL}(a,\mathcal{U})=\log\frac{H}{\delta}/\mathcal{D}_{\min}(a,\mathcal{U})$ samples (i.e., average worst case). This quantity should be compared to the mean number of samples per stage in Theorem 1, which is lower than GJL’s from the maximality of $\tau_{GJL}(a,\mathcal{U})$. Now, we observe that, conditioned on success, if both algorithms use the same action sequences, then both algorithms use the same alive hypotheses at each stage. Consequently, our algorithm uses fewer samples on average than GJL.
Now, we shift our focus to show asymptotic optimality. Accordingly, we bound the error probability of our strategy, given in the following lemma:
@@
For any action $a$, there exists some $\eta_{a}\in(0,1)$ such that (i) $\Delta\mathcal{D}_{ijk}>\eta_{a}$. (ii) The current action can still distinguish between at least two hypotheses. (iii) for any $\varepsilon_{a}\leq\eta_{a}$, the per-stage error probability cannot exceed $\frac{\delta}{H}$. Particularly, $p_{e}\leq\delta$.
Proof:
The detailed proof can be found in Appendix A-A. We first bound the probability for $H_{j}$ to win against $H_{i}$ when $H_{i}$ true but is represented by $H_{k}$ by $\frac{\delta}{H}\times\kappa_{aijk}$ for some $\kappa_{aijk}$ utilizing a similar computation made by Chernoff in [6]. Then, we leverage the clustering to show that there exists some $\eta_{a}$ such that $\kappa_{aijk}\leq 1-\eta_{a}$. Any $\varepsilon_{a}\leq\eta_{a}$ ensures the same, so the claim on the per-stage error probability follows. Once the per-stage error probability bound is established, the result on $p_{e}$ immediately follows from the union bound. ∎
Corollary 1 (Asymptotic Optimality).
For a suitable choice of $\{\varepsilon_{a}\}_{a\in\mathcal{A}}$, Algorithm 1 has vanishing ABR when $\delta\to 0$. Furthermore, the ABR does not exceed $\delta$ when $H\to\infty$.
Proof:
Let $\varepsilon\triangleq\min_{a\in\mathcal{A}}\{\varepsilon_{a}\}$. Combining Lemma 1 and Theorem 1 yields $\frac{\delta}{H^{2}}\mathbb{E}\left[N\right]+p_{e}\leq\frac{\delta}{H^{2}}\times\frac{H}{\varepsilon}\log\frac{H}{\delta}+\delta=\frac{1}{\varepsilon}\times\frac{1}{\frac{H}{\delta}}\log\frac{H}{\delta}+\delta$ , which vanishes when $\delta\to 0$. When $H\to\infty$, only the first addend vanishes, and we are left with $\delta$. ∎
V Numerical Results
In this section, we present numerical results illustrating the performance of our algorithms (with and without clustering) compared to GJL. To this end, we have conducted simulations. In the simulations, we have used uniform priors, i.e., $\pi_{i}=\frac{1}{H}$. For a fair comparison, we picked distributions whose mean can distinguish between different hypotheses, namely, the exponential and unit-variance normal distributions. The number of hypotheses was set to $H=16$, and the DM had $|\mathcal{A}|=17$ actions. The first $H$ actions had a mean of 3 if $H_{a-1}$ is true for $1\leq a\leq H$, and a random mean drawn uniformly from $[0,1]$ otherwise. The last action had the means follow the equation $0.5+0.01\times i$ for $0\leq i\leq H$. All random means were drawn only once. For the clustering algorithm, DBSCAN, we use $\text{minPts}=1$ and (manually tuned) proximity parameter $\varepsilon_{a}=\varepsilon=0.1$ for any $a$.
The simulation results are presented in Figure 2. The dashed lines present the ABR for the case where all actions produce normally distributed samples, whereas the non-dashed lines present the ABR for the case where all actions produce exponentially distributed samples. The orange curves correspond to GJL, and the blue and green curves correspond to our algorithm without and with clustering, respectively. Each algorithm attains vanishing ABR as $\delta\to 0$, but the non-greedy selections made by each instance of our algorithm enjoy significantly lower ABR, attributed to a lower mean number of samples.
[…]
Figure 2: Comparison of the ABR (equation (1)) for $H=16$ using our algorithm (with proximity parameter $\varepsilon\in\{0,0.1\}$) against GJL. The samples obtained are either normally distributed (dashed) or exponentially distributed (non-dashed). The ABR vanishes as $\delta\to 0$ for all instances, but GJL’s greedy selection has a slower decay.
Appendix A Miscellaneous Proofs
A-A Proof of Lemma 1
At first glance, there are two possible error events; the first is that the stage winner (at some stage) is not the representative of the true hypothesis, and the second is the event when two different hypotheses win in the current round. We argue that the latter case never occurs in the following proposition:
Proposition 1.
@@ A-A Proof of Lemma 1
Assume to the contradiction that there are two winners, $i_{r}^{*}$ and $j_{r}^{*}$. Since $i_{r}^{*}$ is a winner, then $L_{i_{r}^{*}j_{r}^{*}}\geq\log\frac{H}{\delta}$. Likewise, $j_{r}^{*}$ is a winner so $L_{j_{r}^{*}i_{r}^{*}}\geq\log\frac{H}{\delta}$. However, due to symmetry, we have $\log\frac{H}{\delta}\leq L_{i_{r}^{*}j_{r}^{*}}=-L_{j_{r}^{*}i_{r}^{*}}\leq-\log\frac{H}{\delta}$, contradiction. ∎
Now that we have established that there are no ties, we are ready to bound the former error probability. Since the number of alive hypotheses can only decrease in each stage, there can be no more than $H$ stages. Consequently, we can compute the error probability of a single stage, and then, by the union bound, the error probability of the algorithm, $p_{e}$, is at most $H$ times the error probability of a single stage. Hence, we will focus on computing the first stage’s error probability for simplicity.